OpenStudy (pokemon23):

A ship left a port and sailed east at the rate of 20mph.One hour later, a second ship left the same port at the rate of 25mph , also traveling east. In how many hours did the second ship overtake the first ship? Show all work?

OpenStudy (pokemon23):

i need someone to show work and explain it ..

OpenStudy (anonymous):

20*1+20x=25x :DD

OpenStudy (anonymous):

now explain what i wrote :P

OpenStudy (pokemon23):

i don't know how you got that answer...

OpenStudy (anonymous):

20 (x+1) = 25 x 20x + 20 = 25x 5x = 20 x = 4 so it's four hours

OpenStudy (anonymous):

moneybird he don't know how to get this equation

OpenStudy (anonymous):

and i don't know how to explain

OpenStudy (anonymous):

First boat has 1hour more than the second boat

OpenStudy (anonymous):

you understand now? :D i think no

OpenStudy (pokemon23):

well he said the train has 1 more than the second boat so its like 20 sailed east at 20mph

OpenStudy (anonymous):

i said same didn't I? :D

OpenStudy (pokemon23):

ya I was playing around with you